need help solving 5radical x+2=155
Answer by algebrahouse.com(1659)

5√(x + 2) = 155
√(x + 2) = 31 {divided both sides by 5}
x + 2 = 961 {squared both sides}
x = 959 {subtracted 2 from both sides}
www.algebrahouse.com